# Márton Benedek 1122    Péter Biró 2211    Gergely Csáji 22 Matthew Johnson 33

### v2

@@ 1 Introduction
We consider partitioned permutation games and IKEPs, so we assume $\ell=\infty$. This assumption is not realistic in kidney exchange, but has also been made for national KEPs to obtain general results of theoretical interest [5, 15, 16, 35] and may have wider applications (e.g. in financial clearing). We also aim to research how stability and the total number of kidney transplants are affected when moving from one extreme ($\ell=2$) to the other ($\ell=\infty$). As such our paper consists of a theoretical part and an experimental part.
We start with our theoretical results (Section 2). Permutation games, i.e. partitioned permutation games of width $1$, have a nonempty core [40], and a core allocation can be found in polynomial time [22]. We generalize these two results to partitioned permutation games of any width $c$, and also show a dichotomy for testing core membership, which is in contrast with the dichotomy for partitioned matching games, where the complexity jump happens at $c=3$ [14].
@@
Our last theoretical result is a randomized XP algorithm with parameter $n$. As we shall prove, derandomizing it requires solving the notorious Exact Perfect Matching problem in polynomial time. The complexity status of the latter problem is still open since its introduction by Papadimitriou and Yannakakis [33] in 1982.
@@
Our theoretical results highlighted severe computational limitations, and we now turn to our simulations. We perform a large-scale experimental study, in which our simulations are strongly guided by our theoretical results. Namely, we note that the algorithm in Theorem 1.5 is not practical for instances of realistic size (which we aim for). Moreover, it is a randomized algorithm. Hence, it it not acceptable either in the setting of kidney exchange, as policy makers would only use solution yielding the maximum number of kidney transplants. Therefore and also due to Theorem 1.4, we formulate the problems of computing a weakly or strongly close optimal solution as integer linear programs, as described in Section 3. This enables us to use an ILP solver. We still exploit the fact that for $\ell=\infty$ (Theorem 1.1) we can find optimal solutions and values $v(S)$ in polynomial time. In this way we can still perform, in Section 4, simulations for IKEPs up to ten countries, so more than the four countries in the simulations for $\ell=3$ [12, 32], but less than the 15 countries in the simulations for $\ell=2$ [10].
@@
For the initial allocations we use two easy-to-compute solution concepts: the benefit value and contribution value and four hard-to-compute solution concepts: the Banzhaf value, Shapley value, nucleolus and tau-value. We also use the modified Banzhaf value called Banzhaf* value from [10], where the credits are incorporated directly into the coalitional values. As in [10], we research both the case where all countries have the same size and the case where countries can be either of small, medium or large size.
@@
Our simulations show, as in [10], that a credit system using strongly close optimal solutions instead of weakly optimal solutions makes an IKEP up to 55% more balanced, without decreasing the overall number of transplants. The exact improvement is determined by the choice of solution concept. Our simulations indicate that the Banzhaf* value yields the best results, namely on average, a deviation of at most 0.90% from the initial allocation (but the differences between the Banzhaf* value with the Banzhaf value and Shapley value are small). Moreover, moving from $\ell=2$ to $\ell=\infty$ yields on average 46% more kidney transplants (using the same simulation instances generated by the data generator [34]). However, the exchange cycles may be very large, in particular in the starting round.
2 Theoretical Results
@@
We first show that finding a core allocation of a partitioned permutation game can be reduced to finding a core allocation of a permutation game. As the latter can be done in polynomial-time [22] (and such a core allocation always exists [40]), the same holds for the former.
Let $(N,v)$ be a partitioned permutation game on a graph $G=(V,A)$ with partition $(V_{1},\ldots,V_{n})$ of $V$. We create a permutation game $(N^{\prime},v^{\prime})$ by splitting each $V_{i}$ into sets of size $1$, i.e., every vertex becomes a player in $N^{\prime}$. Let $x^{\prime}$ be a core allocation of $(N^{\prime},v^{\prime})$. For each set $i\in N$, we set $x(i)=\sum_{v\in V_{i}}x^{\prime}(v)$. It holds that $x(N)=v(N)$, as $(N,v)$ and $(N^{\prime},v^{\prime})$ are defined on the same graph $G$; hence, the weight of a maximum weight cycle packing is unchanged.
@@
Now we show that deciding whether an allocation is in the core can be solved in polynomial time for partitioned permutation games with width $1$, that is, for permutation games. Let $x$ be an allocation. We create a weight function $w_{x}$ over the arcs by setting $w_{x}((u,v))=x(u)-1$. We claim that if there exists a blocking coalition, then there is a blocking coalition that consists of only vertices along a cycle. In order to see this, let $S$ be a blocking coalition, so $x(S)<v(S)$. By definition, $v(S)$ is the maximum size of a cycle packing ${\mathcal{C}}=\{C_{1},C_{2},..,C_{k}\}$ in $G[S]$. For $i=1,\ldots,k$, let $S_{i}$ be the set of vertices in $C_{i}$. From
@@
So we just need to check whether there is a cycle $C$ such that $x(V(C))<|E(C)|=v(V(C))$. By the definition of $w_{x}$, such a cycle exists if and only if $w_{x}$ is not conservative (where conservative means no negative weighted cycles exist). The latter can be decided in polynomial-time, for example with the Bellman-Ford algorithm.
Finally, we show that deciding if an allocation $x$ is in the core of a partitioned permutation game is coNP-complete, even if each $|V_{i}|$ has size $2$ (so $c=2$). Containment in coNP holds, as we can check in polynomial time if a coalition blocks an allocation. To prove hardness, we reduce from a special case of the NP-complete problem Exact $3$-Cover [29, 30].
@@
Finally, we define the allocation $x$, as follows:
  • $x(A_{i})=3-\frac{n+1}{9n^{2}}$ for each $i\in[3n]$,
  • $x(X_{k})=3-\frac{2n-1}{18n^{2}}$ for each $k\in[6n]$, and
  • $x(T_{j}^{l})=1+\frac{1}{9n}$ for each $j\in[3n],l\in[3]$.
@@
Hence, we find that $12h>18n-1+\frac{h}{9n}>18n-$ 1, so $h>9n-1$, but it also cannot be $9n$, as $18n=18n-1+\frac{9n}{9n}$, a contradiction (as there are only $9n$ $T_{j}^{l}$ players).
@@
Hence, in order for $\mathcal{P}$ to block, it must hold that $12h>12n+3m-$ 1, so $h>6n+1.5m-0.5$, contradicting $h\leq 6n+m$, if $m\geq 1$. In the case, when $m=0$, we get that $6n+3h>18n-\frac{2n-1}{3n}+h+\frac{h}{9n}$, so $12h>12n-\frac{2n-1}{3n}+\frac{h}{9n}>12n-$ 1. From this and $h\leq 6n+0$, we get that $h$ must be $6n$. However, then $12n>12n-\frac{2n-1}{3n}+\frac{6n}{9n}>12n$, which is a contradiction again.
@@
If $m\geq 1$, then this implies that $12h>6n+3m-$ 1, so $h>3n+1.5m-0.5$, a contradiction. We conclude that $m=0$. Therefore, $h\leq 3n$ and $12h>6n-\frac{n+1}{3n}+\frac{h}{9n}>6n-$ 1, so $h=3n$.
@@
For instances $(G,{\mathcal{V}},{\mathcal{I}})$, where $G=(V,A)$ is a directed graph, ${\mathcal{V}}=(V_{1},\ldots,V_{n})$ is a partition of $V$ with fixed width $c$, and =${\mathcal{I}}=\{I_{1},\dots,I_{n}\}$ is a set of intervals, the problem of finding a maximum cycle packing of $G$ satisfying ${\mathcal{I}}$ is polynomially solvable if $c=1$, and NP-complete if $c\geq 2$ even if $I_{p}=[1,\infty]$ for every $p\in\{1,\ldots,n\}$ or $I_{p}=[1,1]$ for every $p\in\{1,\ldots,n\}$.
@@
First suppose $c=1$. Let $v_{p}$ be the unique vertex in $V_{p}$. We can assume that each $I_{p}$ contains either 00 or $1$, else no cycle packing satisfying ${\mathcal{I}}$ exists. If $1\notin I_{p}$, we can delete $v_{p}$ and redefine $G$ as the graph that remains. If this decreases the size of the maximum cycle packing, then we conclude that no desired maximum cycle packing exists. Let $U$ be the set of vertices $v_{p}$ for which $0\notin I_{p}$.
The problem reduces to finding a maximum cycle packing such that each vertex in $U$ is covered. For this, we transform $G=(V,A)$ into a bipartite graph $H$ with partition classes $V$ and $V^{\prime}$, where $V^{\prime}$ is a copy of $V$. For each $v\in V\setminus U$ and its copy $v^{\prime}\in V^{\prime}\setminus U^{\prime}$, we add the edge $uu^{\prime}$ with weight 00 (we do not add these for the vertices of $U$). For each $(u,v)\in A$, we add the edge $uv$ with weight $1$. It remains to find in polynomial time a maximum weight perfect matching in $H$, if there is any and check whether its weight is the same as the size of a maximum cycle packing in the original directed graph. If there is a perfect matching with that weight, then in the maximum cycle packing it corresponds to, each vertex in $U$ must be covered with a cycle. In the other direction, if there is a maximum cycle packing covering each vertex in $U$, then the perfect matching it corresponds to has the desired weight and we need no nonexistent $uu^{\prime}$ edge for any $u\in U$ indeed.
@@
  • for each $j\in[3n],l\in[3]$, we have a player $T_{j}^{l}=\{s_{j}^{l},t_{j}^{l}\}$
@@
The interval for each player is $[1,\infty]$. Since the size of the maximum cycle packing of $G$ is $16n$, which is the same as the number of players, if there is a solution that satisfies these intervals, then it also must satisfy the intervals $[1,1]$ for each player. Hence the last two statements of the Lemma are equivalent in this instance.
@@
$\Rightarrow$” First suppose $I$ has an exact $3$-cover $S_{l_{1}},\dots,S_{l_{n}}$. We create a cycle packing $\mathcal{C}$ of $G$. For each $j\in\{l_{1},\dots,l_{n}\}$, we add the cycles $\{(s_{j}^{1},a_{j_{1}}),(a_{j_{1}},s_{j}^{1})\},\{(s_{j}^{2},a_{j_{2}}),(a_{j_{2}},s_{j}^{2})\},\{(s_{j}^{3},a_{j_{3}}),(a_{j_{3}},s_{j}^{3})\}$. For $j\notin\{l_{1},\dots,l_{n}\}$ we add the arcs $(t_{j}^{1},t_{j}^{2}),(t_{j}^{2},t_{j}^{3})$. Finally, for each $i\in[2n]$, we add the arcs $(y_{i},x_{i}),(x_{i},t_{j_{i}}^{1}),(t_{j_{i}}^{3},y_{i})$, where $j_{i}$ is the $i$-th smallest index among the indices $[3n]\setminus\{l_{1},\dots,l_{n}\}$.
@@
Recall that $x_{p}$ denotes the target for the number of arcs $(u,v)$ with $v\in V_{p}$ that belong to some directed cycle of ${\mathcal{C}}$. Letting each $I_{p}=[x_{p},x_{p}]$ and applying Lemma 1, we see that finding a cycle packing where each $s_{p}({\mathcal{C}})$ is equal $x_{p}$ (so differs by at most 00) is NP-complete. Thus it is NP-hard to find the maximum cycle packing that minimizes $d_{1}({\mathcal{C}})=\max_{p\in N}\{|x_{p}-s_{p}({\mathcal{C}})|\}$; that is, to find a solution that is weakly close to a given target and similarly it is also hard to find a strongly close solution. ∎
@@
For $q=2$, this problem is also known as Exact Perfect Matching, which, as mentioned, was introduced by Papadimitriou and Yannakakis [33] and whose complexity status is open for more than 40 years. In the remainder of this section, we will give both a reduction to this problem and a reduction from this problem. We start with the former in the proof of our next result (Theorem 2.1), from which Theorem 1.5 immediately follows.
@@
As each $k_{i}$ can only have a value between 00 and $|E|=|A|+n=O(|A|)$, the above reduction implies that the set of unordered deviation vectors has size $|A|^{O(n)}$ for any allocation $x$ for $(N,v)$. We can find each of these vectors in $|A|^{O(n)}$ time by a randomized algorithm, as $q$-Exact Perfect Matching is solvable in $|E|^{O(q)}$ time with $q$ colours with a randomized algorithm [27]. ∎
@@
Let the first player’s set $V_{1}$ consist of all vertices $v_{e}$, for which $e$ is a red edge in $E$ and let the other player’s vertex set be $V_{2}=V(G)\setminus V_{1}$. This defines a partitioned permutation game $(N,v)$. Let the target allocation $x$ be given by $x=(k_{1},3n-k_{1})$.

### v3

@@ 1 Introduction
We consider partitioned permutation games and IKEPs, so we assume $\ell=\infty$. The latter assumption is not realistic in kidney exchange. However, the assumption has also been made for national KEPs in order to obtain general results of theoretical interest [ACDM21, BKKV23, BMR09, RSU04]. Moreover, results of studies with $\ell=\infty$ may have wider applications, for instance, in portfolio compression in financial clearing [ER2021], time exchange [ACsEE2021], and shift-reallocation [MW2021]. In particular, we wanted to research how the stability of an IKEP and the total number of kidney transplants are affected when moving from one extreme ($\ell=2$) to the other ($\ell=\infty$).
As mentioned, the KEP model we use is the most basic version, which has been used in the literature for conducting computer simulations on generated instance. In particular, this allows us to make a clean comparison with a previous experimental study for $\ell=2$ [BBPY24]. That is, we ignore the diverse sets of complex, hierarchical optimization criteria that are used in European kidney exchange [Bi_etal2019, Bi_etal2021] and only maximize the number of transplants in each matching run. We do note that this objective is always the first objective in every used hierarchy apart from the hierarchy used in the UK, where it is the second objective. We also ignore the presence of non-directed donors and the possibility of having multiple donors registering for one patient. To further justify our choice of not considering non-directed donors, we note that such donors are illegal in some European countries, such as France and Hungary.
Our paper consists of a theoretical part and an experimental part. We start with our theoretical results (Section 2). Permutation games, i.e. partitioned permutation games of width $1$, have a nonempty core [TPPR84], and a core allocation can be found in polynomial time [CT86]. We generalize these two results to partitioned permutation games of any width $c$, and also show a complexity dichotomy for testing core membership, which is in contrast with the complexity dichotomy for partitioned matching games, where the complexity jump happens at $c=3$ [BBKPP].
@@
We refer to Table 1 for an overview of all the known and new results that we discussed above.
Our last theoretical result is a randomized XP algorithm with parameter $n$. As we shall prove, derandomizing it requires solving the notorious Exact Perfect Matching problem in polynomial time. The complexity status of the latter problem is still open since its introduction by Papadimitriou and Yannakakis [PY82] in 1982.
@@
Our theoretical results highlighted severe computational limitations, and we now turn to our simulations. We perform a large-scale experimental study, in which our simulations are strongly guided by our theoretical results. Namely, we note that the algorithm in Theorem 1.5 is not practical for instances of realistic size (which we aim for). Moreover, it is a randomized algorithm. Hence, it it not acceptable either in the setting of kidney exchange, as policy makers would only use solution yielding the maximum number of kidney transplants. Therefore and also due to Theorem 1.4, we formulate the problems of computing a weakly or strongly close optimal solution as integer linear programs, as described in Section 3. This enables us to use an ILP solver. We still exploit the fact that for $\ell=\infty$ (Theorem 1.1) we can find optimal solutions and values $v(S)$ in polynomial time. In this way we can still perform, in Section 4, simulations for IKEPs up to ten countries, so more than the four countries in the simulations for $\ell=3$ [BGKPPV20, KNPV20], but less than the 15 countries in the simulations for $\ell=2$ [BBPY24].
@@
For the initial allocations we use two easy-to-compute solution concepts: the benefit value and contribution value and four hard-to-compute solution concepts: the Banzhaf value, Shapley value, nucleolus and tau-value. In particular, it follows from a well-known alternative definition of the nucleolus [MPS79] that the nucleolus is a core allocation for every game with a nonempty core. As by Theorem 1.3, every partitioned permutation game has a nonempty core, this means that we do not need to consider the core as a separate solution concept in our simulations. Finally, we use the modified Banzhaf value called Banzhaf* value from [BBPY24], where the credits are incorporated directly into the coalitional values. As in [BBPY24], we research both the case where all countries have the same size and the case where countries can be either of small, medium or large size.
@@
Our simulations show, as in [BBPY24], that a credit system using strongly close optimal solutions instead of weakly optimal solutions makes an IKEP up to 56% more balanced [^1] As discussed in Section 4, using the Banzhaf value with $n=10$ and equal country sizes yields an average relative deviation of 1.97% when selecting weakly close optimal solutions and an average relative deviation of 0.86% when selecting strongly close optimal solutions. So, the improvement in balancedness is $(1.97-0.86)/1.97=56\%$. For the other solution concepts, the improvements are between 36% and 54%. without decreasing the overall number of transplants. Our simulations also indicate that out of the seven solution concepts, using the Banzhaf* value leads to the smallest deviations, namely on average, a deviation of at most 0.90% from the initial allocation (but the differences between the Banzhaf* value and the Banzhaf value or the Shapley value are small). Moreover, moving from $\ell=2$ to $\ell=\infty$ yields on average 46% more kidney transplants (using the same simulation instances generated by the data generator [PT21]).
For our simulations, we also kept track of the length of the exchange cycles. It turned out that these may be very large, in particular in the starting round. Previously, the relation between increase in transplants versus increase in cycle length was also researched by Biró [BMR09], namely for $\ell=2$, $\ell=3$, and $\ell=\infty$. However, their simulations were done in the context of the UK KEP only.
$\ell=2$ $\ell\in\{3,4,\ldots\}\;\;$ $\ell=\infty$
finding an optimal solution poly [ABS07] NP-h [ABS07] poly [ABS07]
testing core nonemptyiness poly if $c\leq 2$ NP-h poly [TPPR84]
coNP-hard if $c\geq 3$ [BBKPP]
finding a core allocation poly if $c\leq 2$ NP-h poly [CT86]
coNP-h if $c\geq 3$ [BBKPP]
deciding core membership poly if $c\leq 2$ NP-h poly if $c=1$
coNP-c if $c\geq 3$ [BBKPP] coNP-c if $c\geq 2$
finding a weakly close solution to target allocation poly [BBKPP] NP-h NP-h
finding a strongly close solution to target allocation poly [BBKPP] NP-h NP-h
Table 1: Survey of complexity results relevant for IKEPs, for fixed cycle length $\ell=2$, $\ell\in\{3,4,\ldots\}$ and $\ell=\infty$, respectively, where $c$ is the width of the associated partitioned $\ell$-permutation game $(N,v)$. Here, poly stands for polynomial-time solvable; coNP-c for coNP-complete; NP-h for NP-hard; and coNP-h means coNP-hard. Results in purple are new results shown in this paper, whereas non-referenced results follow directly from the referenced results.
2 The Proofs of Our Theoretical Results
@@
We first show that finding a core allocation of a partitioned permutation game can be reduced to finding a core allocation of a permutation game. As the latter can be done in polynomial-time [CT86] (and such a core allocation always exists [TPPR84]), the same holds for the former.
Let $(N,v)$ be a partitioned permutation game on a graph $G=(V,A)$ with partition $(V_{1},\ldots,V_{n})$ of $V$. We create a permutation game $(N^{\prime},v^{\prime})$ by splitting each $V_{i}$ into sets of size $1$, i.e., every vertex becomes a player in $N^{\prime}$. Let $x^{\prime}$ be a core allocation of $(N^{\prime},v^{\prime})$. For each player $i\in N$, we set ${x_{i}}=\sum_{v\in V_{i}}{x^{\prime}_{v}}$, that is, we sum the $x^{\prime}$ values of the players in $N^{\prime}$ corresponding to the vertices that $i$ controls in $G$. It holds that $x(N)=v(N)$, as $(N,v)$ and $(N^{\prime},v^{\prime})$ are defined on the same graph $G$; hence, the weight of a maximum weight cycle packing is unchanged.
@@
Now we show that deciding whether an allocation is in the core can be solved in polynomial time for partitioned permutation games with width $1$, that is, for permutation games. Let $x\in\mathbb{R}^{N}$ be an allocation. We create a weight function $w_{x}$ over the arcs by setting $w_{x}((u,v))={x_{u}}-1$. We claim that if there exists a blocking coalition, then there is a blocking coalition that consists of only vertices along a cycle. In order to see this, let $S$ be a blocking coalition, so $x(S)<v(S)$. By definition, $v(S)$ is the maximum size of a cycle packing ${\mathcal{C}}=\{C_{1},C_{2},..,C_{k}\}$ in $G[S]$. For $i=1,\ldots,k$, let $S_{i}$ be the set of vertices in $C_{i}$. From
@@
So we just need to check whether there is a cycle $C_{i}$ with vertex set $S_{i}$ such that $x(S_{i})<|S_{i}|$. Note that $\sum_{(u,v)\in C_{i}}w_{x}(u,v)=x(S_{i})-|S_{i}|$. Hence, such a cycle exists if and only if $w_{x}$ is not conservative (where conservative means no negative weighted cycles exist). The latter can be decided in polynomial-time, for example with the Bellman-Ford algorithm.
Finally, we show that deciding if an allocation $x$ is in the core of a partitioned permutation game is coNP-complete, even if each $|V_{i}|$ has size $2$ (so $c=2$). Containment in coNP holds, as we can check in polynomial time if a coalition blocks an allocation. To prove hardness, we reduce from a special case of the NP-complete problem Exact $3$-Cover [GJWZ96, HDRB08].
@@
Finally, we define the allocation $x{\in\mathbb{R}^{N}}$, as follows:
  • ${x_{A_{i}}}=3-\frac{n+1}{9n^{2}}$ for each $i\in[3n]$,
  • ${x_{X_{k}}}=3-\frac{2n-1}{18n^{2}}$ for each $k\in[6n]$, and
  • ${x_{T_{j}^{l}}}=1+\frac{1}{9n}$ for each $j\in[3n],l\in[3]$.
@@
Hence, we find that $2h>18n-1+\frac{h}{9n}>18n-1$, so h>9n1h>9n-1, but it also cannot be $9n$, as $18n=18n-1+\frac{9n}{9n}$, a contradiction (as there are only $9n$ $T_{j}^{l}$ players).
@@
Hence, in order for $\mathcal{P}$ to block, it must hold that 2h>12n+3m12h>12n+3m-1, so h>6n+1.5m0.5h>6n+1.5m-0.5, contradicting $h\leq 6n+m$, if $m\geq 1$. In the case, when $m=0$, we get that $6n+3h>18n-\frac{2n-1}{3n}+h+\frac{h}{9n}$, so $2h>12n-\frac{2n-1}{3n}+\frac{h}{9n}>12n-1$. From this and $h\leq 6n+0$, we get that $h$ must be $6n$. However, then $12n>12n-\frac{2n-1}{3n}+\frac{6n}{9n}>12n$, which is a contradiction again.
@@
If $m\geq 1$, then this implies that 2h>6n+3m12h>6n+3m-1, so h>3n+1.5m0.5h>3n+1.5m-0.5, a contradiction. We conclude that $m=0$. Therefore, $h\leq 3n$ and $2h>6n-\frac{n+1}{3n}+\frac{h}{9n}>6n-1$, so $h=3n$.
@@
For instances $(G,{\mathcal{V}},{\mathcal{I}})$, where $G=(V,A)$ is a directed graph, ${\mathcal{V}}=(V_{1},\ldots,V_{n})$ is a partition of $V$ with fixed width $c$, and ${\mathcal{I}}=\{I_{1},\dots,I_{n}\}$ is a set of intervals, the problem of finding a maximum cycle packing of $G$ satisfying $\mathcal{I}$} is polynomially solvable if $c=1$, and NP-complete if $c\geq 2$ even if $I_{p}=[1,\infty]$ for every $p\in\{1,\ldots,n\}$ or $I_{p}=[1,1]$ for every $p\in\{1,\ldots,n\}$.
@@
First suppose $c=1$. Let $v_{p}$ be the unique vertex in $V_{p}$. We can assume that each $I_{p}$ contains either 0 or $1$, else no cycle packing satisfying $\mathcal{I}$} exists. If $1\notin I_{p}$, we can delete $v_{p}$ and redefine $G$ as the graph that remains. If this decreases the size of the maximum cycle packing, then we conclude that no desired maximum cycle packing exists. Let $U$ be the set of vertices $v_{p}$ for which $0\notin I_{p}$.
The problem reduces to finding a maximum cycle packing such that each vertex in $U$ is covered. For this, we transform $G=(V,A)$ into a bipartite graph $H$ with partition classes $V$ and $V^{\prime}$, where $V^{\prime}$ is a copy of $V$. For each $v\in V\setminus U$ and its copy $v^{\prime}\in V^{\prime}\setminus U^{\prime}$, we add the edge $uu^{\prime}$ with weight 0 (we do not add these for the vertices of $U$). For each $(u,v)\in A$, we add the edge $uv$ with weight $1$. It remains to find in polynomial time a maximum weight perfect matching in $H$, if there is any and check whether its weight is the same as the size of a maximum cycle packing in the original directed graph. If there is a perfect matching with that weight, then in the maximum cycle packing it corresponds to, each vertex in $U$ must be covered with a cycle. In the other direction, if there is a maximum cycle packing covering each vertex in $U$, then the perfect matching it corresponds to has the desired weight and we need no nonexistent $uu^{\prime}$ edge for any $u\in U$ indeed.
@@
  • for each $j\in[3n],l\in[3]$, we have a player $T_{j}^{l}=\{s_{j}^{l},t_{j}^{l}\}$ .
@@
The interval for each player is $1,\infty]$. Since the size of the maximum cycle packing of $G$ is $16n$, which is the same as the number of players, if there is a solution that satisfies these intervals, then it also must satisfy the intervals $[1,1]$ for each player. Hence the last two statements of the Lemma are equivalent in this instance.
@@
$\Rightarrow$” First suppose $I$ has an exact $3$-cover $S_{l_{1}},\dots,S_{l_{n}}$. We create a cycle packing $\mathcal{C}$ of $G$. For each $j\in\{l_{1},\dots,l_{n}\}$, we add the cycles $\{(s_{j}^{1},a_{j_{1}}),(a_{j_{1}},s_{j}^{1})\},\{(s_{j}^{2},a_{j_{2}}),(a_{j_{2}},s_{j}^{2})\},\{(s_{j}^{3},a_{j_{3}}),(a_{j_{3}},s_{j}^{3})\}$. For $j\notin\{l_{1},\dots,l_{n}\}$ we add the arcs $(t_{j}^{1},t_{j}^{2}),(t_{j}^{2},t_{j}^{3})$. Finally, for each $i\in[2n]$, we add the arcs $(y_{i},x_{i}),(x_{i},t_{j_{i}}^{1}),(t_{j_{i}}^{3},y_{i})$, where $j_{i}$ is the $i$-th smallest index among the indices $3n]\setminus\{l_{1},\dots,l_{n}\}$.
@@
Recall that $x_{p}$ denotes the target for the number of arcs $(u,v)$ with $v\in V_{p}$ that belong to some directed cycle of $\mathcal{C}$}. Letting each $I_{p}=[x_{p},x_{p}]$ and applying Lemma 1, we see that finding a cycle packing where each $s_{p}({\mathcal{C}})$ is equal $x_{p}$ (so differs by at most 0) is NP-complete. Thus it is NP-hard to find the maximum cycle packing that minimizes $d_{1}({\mathcal{C}})=\max_{p\in N}\{|x_{p}-s_{p}({\mathcal{C}})|\}$; that is, to find a solution that is weakly close to a given target and similarly it is also hard to find a strongly close solution. ∎
@@
For $q=2$, this problem is also known as Exact Perfect Matching, which, as mentioned, was introduced by Papadimitriou and Yannakakis [PY82] and whose complexity status is open for more than 40 years. In the remainder of this section, we will give both a reduction to this problem and a reduction from this problem. We start with the former in the proof of our next result (Theorem 2.1), from which Theorem 1.5 immediately follows.
@@
As each $k_{i}$ can only have a value between 0 and $|E|=|A|+n=O(|A|)$, the above reduction implies that the set of unordered deviation vectors has size $|A|^{O(n)}$ for any allocation $x$ for $(N,v)$. We can find each of these vectors in $|A|^{O(n)}$ time by a randomized algorithm, as $q$-Exact Perfect Matching is solvable in $|E|^{O(q)}$ time with $q$ colours with a randomized algorithm [GKMST2012]. ∎
@@
Let the first player’s set $V_{1}$ consist of all vertices $v_{e}$, for which $e$ is a red edge in $E$ and let the other player’s vertex set be $V_{2}=V\setminus V_{1}$. This defines a partitioned permutation game $(N,v)$. Let the target allocation $x$ be given by $x=(k_{1},3n-k_{1})$.